# Winner-takes-all for Multivariate Probabilistic Time Series Forecasting

### v1

@@ C.3.2 Baselines
Elec. 315611 $\pm$ 1609 451750 $\pm$ 52854 276601 $\pm$ 19341 371934 $\pm$ 16589 4584308 $\pm$ 84474 433494 $\pm$ 23503 220372 $\pm$ 29409 245904 $\pm$ 26410
Exch. 0.426 $\pm$ 0.0045 0.633 $\pm$ 0.0662 0.225 $\pm$ 0.0176 0.606 $\pm$ 0.016 2.075 $\pm$ 0.0376 1.156 $\pm$ 0.155 0.031 $\pm$ 0.0065 0.042 $\pm$ 0.0123
Solar 5899 $\pm$ 15 3924 $\pm$ 330 4249 $\pm$ 240 6505 $\pm$ 1897 10163 $\pm$ 198 3637 $\pm$ 354 3391 $\pm$ 820 2195 $\pm$ 257
Traffic 18.782 $\pm$ 0.014 20.755 $\pm$ 0.753 10.89 $\pm$ 1.079 12.9 $\pm$ 1.391 69.662 $\pm$ 2.345 17.644 $\pm$ 0.234 5.766 $\pm$ 0.165 5.715 $\pm$ 0.245
Taxi 4385.41 $\pm$ 9.92 5330.4 $\pm$ 416.49 5262.07 $\pm$ 592.83 3728.29 $\pm$ 6.39 7713.98 $\pm$ 569.13 4104.29 $\pm$ 361.39 712.56 $\pm$ 291.79 703.29 $\pm$ 171.0
Wiki. 19692927 $\pm$ 233519 18818790 $\pm$ 1248330 2650865 $\pm$ 422885 2518198 $\pm$ 233471 9925416 $\pm$ 276424 12729658 $\pm$ 367297 9530 $\pm$ 9256 271611 $\pm$ 311180
@@ C.4 Evaluation on financial data
FLOPs 3.89 ×$\times 10^{7}$ 9.81 ×$\times 10^{7}$ 9.13 ×$\times 10^{8}$ 1.74 ×$\times\mathbf{10^{5}}$ 1.94 ×$\times 10^{7}$ 9.98 ×$\times 10^{5}$
@@
Setup. We trained and evaluated the previously introduced neural-based baselines on this dataset, using $K=4$ hypotheses. Each model was trained for $100$ epochs, with $100$ iterations per epoch and a batch size of $64$. The prediction length and context length were both set to $24$. The experimental setup was the same as in the previous benchmark, except that we applied Z-score normalization instead of mean scaling during training, which better handles the wide range of asset price scales. With the exception of Tactis2, which already uses Z-score normalization, we observed that this change significantly improved the performance of all baselines, except for DeepAR where we retained mean scaling.
Results. Quantitative results on quality (Distortion, CRPS, RMSE), smoothness (Total Variation), and computational cost (FLOPs) are presented in Table 12. TimeMCL produces the smoothest predictions overall. Among non-transformer-based baselines (TimeGrad, DeepAR, and TempFlow), TimeMCL consistently outperforms the others in terms of quality. In this setup, it also performs competitively with Tactis2, which was the strongest competitor in the previous benchmark. Unlike in the previous benchmark, we found that the Transformer-based variant of TempFlow (Trf.TempFlow) now outperforms the original TempFlow and slightly surpasses TimeMCL in quality metrics, though it requires nearly $40$ times more FLOPs. In the future, we plan to refine this comparison by implementing a Transformer-based version of TimeMCL.
Cryptocurrency price predictions for 04/01/2025 are visualized in Figure 6. Each row represents a target dimension, and each column corresponds to a method listed in Table 12. Following the notation introduced in Figure, TimeMCL predictions are shown in shades of blue, with color intensity reflecting the associated score. We observe that TimeMCL produces smoother predictions compared to other methods and effectively captures rare modes in the conditional distribution, as illustrated in the last row. For clarity, only a subset of the cryptocurrencies is shown in the figure; however, all models were trained to jointly predict all cryptocurrencies.

### v2

@@ C.3.2 Baselines
Elec. 315611 $\pm$ 1858 451750 $\pm$ 61031 284232 $\pm$ 23269 372443 $\pm$ 18499 4584308 $\pm$ 97542 434097 $\pm$ 34623 220375 $\pm$ 33961 245908 $\pm$ 30505
Exch. 0.426 $\pm$ 0.0052 0.633 $\pm$ 0.0764 0.237 $\pm$ 0.0355 0.606 $\pm$ 0.0208 2.075 $\pm$ 0.0434 1.104 $\pm$ 0.1582 0.031 $\pm$ 0.0075 0.042 $\pm$ 0.0141
Solar 5899 $\pm$ 17 3924 $\pm$ 381 4421 $\pm$ 828 5383 $\pm$ 2173 10163 $\pm$ 229 3653 $\pm$ 488 3391 $\pm$ 947 2195 $\pm$ 297
Traffic 18.782 $\pm$ 0.016 20.56 $\pm$ 0.932 11.063 $\pm$ 0.874 12.991 $\pm$ 1.661 69.603 $\pm$ 2.61 18.114 $\pm$ 0.512 5.766 $\pm$ 0.19 5.714 $\pm$ 0.283
Taxi 4385.41 $\pm$ 11.46 5331.4 $\pm$ 517.54 5452.77 $\pm$ 324.07 4232.16 $\pm$ 713.9 7932.26 $\pm$ 711.87 4147.2 $\pm$ 362.87 709.86 $\pm$ 332.07 701.61 $\pm$ 199.35
Wiki 19692927 $\pm$ 269644 18843511 $\pm$ 1553022 2634547 $\pm$ 597323 2458503 $\pm$ 151593 9983566 $\pm$ 434670 12410909 $\pm$ 319517 9532 $\pm$ 10690 271611 $\pm$ 359320
@@ C.4 Evaluation on financial data
FLOPs 3.89 $\times 10^{7}$ 9.81 $\times 10^{7}$ 9.13 $\times 10^{8}$ 1.74 $\times\mathbf{10^{5}}$ 1.94 $\times 10^{7}$ 9.98 $\times 10^{5}$
@@
Setup. We trained and evaluated the previously introduced neural-based baselines on this dataset, using $K=4$ hypotheses. Each model was trained for $100$ epochs, with $100$ iterations per epoch and a batch size of $64$. The prediction length and context length were both set to $24$. The experimental setup was the same as in the previous benchmark, except that we applied Z-score normalization instead of mean scaling during training, which better handles the wide range of asset price scales. With the exception of Tactis2, which already uses Z-score normalization, we observed that this change significantly improved the performance of all baselines, except for DeepAR where we retained mean scaling. Note that, because the validation losses were monotonically decreasing, we used the final training checkpoint for testing.
Results. Quantitative results on quality (Distortion, CRPS, RMSE), smoothness (Total Variation), and computational cost (FLOPs) are presented in Table 12. TimeMCL produces the smoothest predictions overall. Among non-transformer-based baselines (TimeGrad, DeepAR, and TempFlow), TimeMCL consistently outperforms the others in terms of quality. In this setup, it also performs competitively with Tactis2, which was a strong competitor in the previous benchmark. Unlike in the previous benchmark, we found that the Transformer-based variant of TempFlow (Trf.TempFlow) now outperforms the original TempFlow and slightly surpasses TimeMCL in quality metrics, though it requires nearly $40$ times more FLOPs. In the future, we plan to refine this comparison by implementing a Transformer-based version of TimeMCL.
Cryptocurrency price predictions are visualized in Figure 6. Each row represents a target dimension, and each column corresponds to a method listed in Table 12. Following the notation introduced in Figure, TimeMCL predictions are shown in shades of blue, with color intensity reflecting the associated score. We observe that TimeMCL produces smoother predictions compared to other methods and effectively captures different modes in the conditional distribution. For clarity, only a subset of the cryptocurrencies is shown in the figure; however, all models were trained to jointly predict all cryptocurrencies.